# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v9

@@ 6. Illustrative Examples
7. Significance and Applications to Physics and STEM Disciplines
Connections with torsion and product structures appear naturally in a wide range of mathematical and physical models. In string theory, supergravity, and related frameworks, metric connections with totally skew-symmetric torsion play a central role in flux compactifications, generalized geometries, and effective field theories on product manifolds such as $M_{4}\times X$. In these settings, the holonomy of the canonical connection governs the coupling between the external (space-time) and internal (compactified) factors, and irreducibility prevents the decoupling of physical degrees of freedom.
Our lower bound provides an efficient mechanism for establishing such irreducibility directly from global invariants - particularly the mixed tensor rank and the corrected quantity $r^{\sharp}$ - without computing curvature tensors or higher covariant derivatives. This is especially relevant in high-dimensional backgrounds where explicit curvature calculations are notoriously difficult or mathematically intractable.
A similar phenomenon arises in gauge theory, condensed matter, and materials science, where product manifolds model layered, composite, or multi-component systems. The holonomy algebra encodes the geometric interactions between these components; detecting whether such interactions are essential (irreducible) or merely apparent becomes a fundamental structural question. The criterion developed in this paper identifies irreducibility based solely on the existence and non-parallelity of mixed differential forms, providing a geometric diagnostic that is robust under perturbations of the underlying metric and torsion.
Applications also extend to geometric analysis and computational modeling. In multi-scale PDE systems defined on product domain (such as transport-diffusion equations, elasticity models, or geometric flows) the holonomy of a natural geometric connection determines how different subsystems interact and whether the governing equations admit a decoupled formulation. Our lower bound offers a rigorous obstruction to such decoupling, enabling researchers to understand when full coupling is intrinsic to the geometry. This is particularly useful in numerical simulations, where direct computation of curvatures is expensive and unstable, while global invariants such as parallel-form spaces are straightforward to detect algorithmically.
Overall, the ability to certify irreducible holonomy through topological or geometric invariants provides a powerful tool across physics, geometry, and applied sciences. It eliminates the need for explicit curvature computations and transforms a classically analytic problem into a global geometric one, opening the door to applications in any setting where torsion, product structures, and holonomy play a foundational role.
8. Concluding Remarks and Perspectives
We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with 3-form torsion representing a mixed cohomology class via its harmonic projection $\omega_{h}$, the off-diagonal holonomy dimension admits a quantified lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K},$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\tilde{\Psi}_{p}$ is the sum of the kernels of the linear maps $\Psi_{p},\tilde{\Psi}_{p}$ that encode the obstruction to off-diagonal curvature. The resulting lower bound is metric-invariant modulo parallel-form preserving deformations, thus providing the first computable obstruction to reducible holonomy on product manifolds. This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace contained within the mixed class. The examples above illustrate the computation of $r^{\sharp}$ in simple geometric settings. When the torsion is purely harmonic and the parallel-form spaces are minimal, the bound is often saturated in explicit constructions (e.g., on flat tori with parallel torsion). It is natural to conjecture that equality holds generically for purely harmonic torsion, while strict inequality may occur when the coexact or exact parts of $T$ generate additional off-diagonal curvature directions, or when the parallel-form spaces are larger. A systematic classification of the cases where the bound is sharp remains an interesting direction for future work.
@@ 8. Concluding Remarks and Perspectives
A further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].

### v10

@@ 6. Illustrative Examples
7. Conclusion
We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$—specifically, products of compact oriented surfaces or four‑dimensional $(3,1)$-splittings with rank‑one mixed classes—and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\;\geq\;r^{\sharp}=r-\dim\mathcal{K},$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\operatorname{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the sum of the kernels of the linear obstruction maps. The bound is metric‑invariant under deformations preserving the parallel‑form strata $\mathcal{P}_{k}(M_{i})$. This provides the first computable obstruction to reducible holonomy on product manifolds within the admissible class.
The examples in Section 6 illustrate both non‑trivial and trivial cases, while Remark 2.4 shows by explicit construction that the bound fails to be well‑defined without the admissibility hypothesis, establishing sharpness of our assumptions.
When the torsion is purely harmonic and the parallel‑form strata are minimal, the bound is often saturated (e.g., on flat tori with parallel torsion). It is natural to conjecture that equality holds generically in this setting, while strict inequality may occur when the coexact/exact parts of $T$ generate additional off‑diagonal curvature or when the parallel‑form spaces are larger. A systematic classification of sharp cases remains an interesting open problem.
A further direction is to explore extensions beyond the admissible setting. Imposing Kruskal‑rank constraints on harmonic representatives could yield intrinsic mixed factor spaces even when $b_{2}(M_{i})>1$, potentially treating Calabi–Yau threefolds or $G_{2}\times S^{1}$ backgrounds. Such extensions would require new algebraic tools and are left for future investigation.
@@ 7. Conclusion
Another further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].